Sid's Skins makes a variety of covers for electronic organizers and portable music players. The company's designers have discovered a market for a new clear plastic covering with college logos for a popular music player. Market research indicates that a cover like this would sell well in the market priced at $20.50. Sid's desires an operating profit of 25 percent of costs.

Required:

What is the highest acceptable manufacturing cost for which Sid's would be willing to produce the cover? (Round your answer to 2 decimal places.)

Highest Acceptable Manufacturing Costs:
Answer #1


Solution: Operating profit=Sales-cost let the cost be a 0.25a-20.50-a 0.25a+a=20.50 1.25a=20.50 a=20.50/1.25 a=Highest Acce